In a chemistry class, 7 liters of a 4% silver iodide solution must be mixed with a 10% solution to get a 6% solution. How many liters of a 10% solution are needed?
A. 4.5L
B. 2.5L
C. 7.0L
D. 3.5L

Answer by Paul(988)

14x+6(7-x)=10(7)
14x+42-6x=70
8x=28
x=3.5
Hence, 3.5L of 10% solution is needed.